# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

### v1

@@ Abstract
The analysis of Stochastic Gradient Descent (SGD) often relies on making some assumption on the variance of the stochastic gradients, which is usually not satisfied or difficult to verify in practice. This paper contributes to a recent line of works which attempt to provide guarantees without making any variance assumption, leveraging only the (strong) convexity and smoothness of the loss functions. In this context, we prove new theoretical bounds derived from the monotonicity of a simple Lyapunov energy, improving the current state-of-the-art and extending their validity to larger step-sizes. Our theoretical analysis is backed by a Performance Estimation Problem analysis, which allows us to claim that, empirically, the bias term in our bounds is tight within our framework.
@@ 1 Introduction
Given $n\geq 1$ convex and smooth real-valued functions $f_{1},\dots,f_{n}\colon\mathbb{R}^{d}\to\mathbb{R}$, we consider the problem given by
$\min_{x\in\mathbb{R}^{d}}f(x),\quad\text{where }f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x).$ (Finite-Sum)
We denote $S=\operatorname{argmin}(f)$, and assume it to be nonempty.
The Stochastic Gradient Descent (SGD) algorithm (robbins_stochastic_1951) is one of the most popular approaches for solving Problem (Finite-Sum), and is widely used for large-scale machine learning and stochastic optimization problems. In its standard version, at each iteration, SGD performs a gradient step using one of the functions in the sum, which is chosen at random. More precisely, given the iterate $x_{t}$, SGD picks $i_{t}\in\{1,\dots,n\}$ uniformly at random, and computes $x_{t+1}$ via
@@
where $\gamma>0$ is the step-size. Intuitively, SGD is able to progress towards the solutions of the problem because it follows a direction that, on average, is an unbiased estimator for $\nabla f$. The simplicity and efficiency of SGD make it the go-to algorithm for training deep neural networks and other models on massive datasets, where computing the entire gradient $\nabla f$ is intractable or too expensive.
Complexity of SGD. Quantitative results on the behavior of SGD consist in upper bounds for a metric $\Delta(T)$, which is to be interpreted as an optimality gap at iteration $T$. Typical choices include the difference between the expected value of $f$ at a point obtained after $T$ iterations and its minimum value, or the expected square-distance from such a point to a solution. In most known results, these upper bounds are expressed as the sum of two terms, taking the form
@@
where the bias term typically vanishes when $T\to+\infty$, and where the variance term can ideally be made arbitrarily small by taking a small enough step-size. Ideally, we would like the bias term to be equal to the rate of convergence of Gradient Descent (GD).
Assumptions for SGD. Classical studies on SGD typically assume that each function $f_{i}$ is $L$-smooth and either convex, strongly convex or satisfies a Polyak-Lojasiewicz inequality. Some works assume relaxations of these hypotheses, making assumptions under the form of variational inequalities, such as the expected smoothness (gower_sgd_2019), the $(L_{0},L_{1})$-smoothness (zhang_why_2020), the expected residual (gower_stochastic_2021), or the ABC condition (khaled_better_2023). Other works assume that only the function $f$ is smooth or convex. For the sake of readability, we will assume that each of the functions $f_{i}$ is convex and smooth, even though we show in the Appendix A that we simply need an assumption analog to expected smoothness.
Besides the assumptions on the problem itself, it is quite standard to make an assumption on the algorithm, through assumptions on the variance of the stochastic gradients $\nabla f_{i_{t}}(x_{t})$. Many different variance assumptions have been studied in the context of SGD. The oldest and possibly the best-known of such assumptions is the uniform boundedness of the variance (robbins_stochastic_1951; nemirovski_robust_2009; rakhlin_making_2012; schmidt_fast_2013), or a variant, namely:
$\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)-\nabla f(x)\|^{2}}\right]\mathclose{}<+\infty\quad\text{or}\quad\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)\|^{2}}\right]\mathclose{}<+\infty.$
The latter assumption strictly implies the former, and is usually made in non-smooth stochastic optimization (bubeck_convex_2015, Section 6), or in online optimization (hazan_introduction_2016), whereas the former is more common in our convex and smooth setting. These assumptions are however unrealistic in practice (bottou_optimization_2018; nguyen_sgd_2018), and are never verified for strongly convex functions. The bounded variance assumption has been relaxed into weaker assumptions, such as the Blum-Gladyshev assumption (blum_approximation_1954; gladyshev_stochastic_1965; alacaoglu_towards_2025), maximal strong growth (schmidt_fast_2013), strong and weak growth (vaswani_fast_2019), or relaxed growth (bertsekas_gradient_2000; bottou_optimization_2018), to name a few. But the problem remains that those assumptions cannot be verified a priori in practice for smooth problems.
In this paper, we make no boundedness assumptions on the variance, a setting already pioneered in (bach_non-asymptotic_2011; needell_stochastic_2016; gower_sgd_2019).
Our contribution. We establish new and improved quantitative results on the complexity of SGD in the convex and $L$-smooth setting, without making any variance assumption. Instead our bounds will rely on the so-called solution variance constant, which is always finite and defined as
$\sigma_{*}^{2}=\frac{1}{m}\sum_{i=1}^{n}\|\nabla f_{i}(x_{*})\|^{2},$
where $x_{*}$ is any minimizer of $f$. Indeed, the constant does not depend on the choice of $x_{*}$, see (garrigos_handbook_2024, Lemma 4.17). Our results take the form
@@
and will be valid for the whole range of step-sizes $\gamma L\in(0,2)$. We focus on obtaining the smallest bias term possible, and on top of improving it with respect to the existing results, we conducted numerical experiments illustrating the sharpness of our bound. To the best of our knowledge this is the first work where a unified and comprehensive study of SGD, without bounded variance assumptions and in the whole range of step-sizes, has been carried out in the convex and smooth setting.
In the convex case, we establish bounds for the expected function value gap $\Delta(T)=\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}$, where $\bar{x}_{T}$ is the Cesàro average of the iterates generated by SGD up to iteration $T$. Our bounds, with simplified notation for ease of presentation, are:
${\rm Bias}(T)=\begin{cases}\tfrac{1}{2\gamma T}&\text{ if }\gamma L\in(0,1),\\ \tfrac{1}{(2-\varepsilon)\gamma T}&\text{ if }\gamma L=1,\varepsilon>0,\\ \tfrac{1}{2\gamma(2-\gamma L)T}&\text{ if }\gamma L\in(1,2),\end{cases}\quad{\rm Variance}(T)=\begin{cases}\tfrac{\gamma}{2(1-\gamma L)}&\text{ if }\gamma L\in(0,1),\\ \tfrac{\gamma(2+\varepsilon)}{\varepsilon(2-\varepsilon)}&\text{ if }\gamma L=1,\varepsilon>0,\\ \tfrac{\exp(T)}{2-\gamma L}&\text{ if }\gamma L\in(1,2).\end{cases}$
In our setting where no assumptions on the variance are made, the only existing bounds are available for step-sizes limited to $\gamma L\in(0,\tfrac{1}{2})$, and provide worse constants (gower_sgd_2021). Our results are the best so far and the first one valid in the whole interval $(0,1)$. Moreover, it matches the bias term previously obtained under the bounded variance assumption (taylor_stochastic_2019). For $\gamma L=1$ we are able to have a bias term as close as possible from the optimal bias $\tfrac{1}{2\gamma L}$, but cannot have equality with a finite variance term. Numerical results suggest that it is not possible to achieve this optimal bias. For $\gamma L\in(1,2)$, which is a completely uncharted territory for this problem, we find that the variance term grows exponentially with $T$, making it impractical. Again, our numerical experiments suggest that one cannot avoid this growth in the variance term. Note that those problems encountered for large step-sizes $\gamma L\in[1,2)$ are typically avoided when making a uniformly bounded variance assumption, see e.g. taylor_stochastic_2019.
@@
As a by-product, our new bounds for SGD with the step-size $\gamma L=1$ makes it possible to study the stochastic proximal algorithm through the lens of SGD. To the best of our knowledge, we establish the first complexity guarantees for this method in a general convex nonsmooth setting.
In the case where each function $f_{i}$ is $\mu$-strongly convex, we provide bounds over the expected distance to the solution, namely $\Delta(T)=\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}$. Our results show that we can take ${\rm Bias}(T)=\phi^{2T}$ where $\phi$ corresponds to the standard rate of GD:
$\phi=\max\{1-\gamma\mu;\gamma L-1\},$
with the exception of the optimal step-size $\gamma=\tfrac{2}{\mu+L}$, for which the bias term must be relaxed to avoid a blow up in the variance. These results improve on (bach_non-asymptotic_2011; needell_stochastic_2016; gower_sgd_2019) by allowing for the full range of step-sizes $\gamma L\in(0,2)$, and by having a sharp bias term. Our numerical experiments suggest again that we cannot improve those results.
@@
The strategy. For each fixed time horizon $T$, we propose a time-dependent energy random sequence
@@
where $\rho$, $(a_{t})$ and $(e_{t})$ are nonnegative. The first ingredient of our Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which typically decreases for deterministic gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. The standard Lyapunov for gradient descent usually contains a term $t(f(x_{t})-\min f)$, which we have replaced here with the sum of the past function gaps, which is of the same order in time. This choice is critical to obtain bounds without variance assumptions (compare with the Lyapunov energies considered by taylor_stochastic_2019). The last term is a negative cumulated sum, where the $e_{t}$’s play the role of a variance term. It is here to compensate the fluctuations caused by the variance in the SGD algorithm, and will allow the Lyapunov energy to decrease.
@@
A decrease of $\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ along the iterations means that $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, which translates into a bound for SGD. Our strategy is therefore to find admissible Lyapunov parameters $\rho$, $(a_{t})$ and $(e_{t})$, which are those that make the energy decrease in expectation. Among those admissible parameters, our goal is to find the ones minimizing, whenever possible, the bias term in our bound.
To this end, we rely on the approach followed in taylor_stochastic_2019 to design Lyapunov potentials in a systematic fashion. It is based on the Performance Estimation Problem (PEP) methodology, which originally appeared in drori_performance_2014, and was further developed in taylor_smooth_2017. A similar line of thought was followed in (fercoq_defining_2024). Automated Lyapunov analyses have also been developed in (taylor_lyapunov_2018; upadhyaya_automated_2025) by combining the PEP approach with the Integral Quadratic Constraints framework (lessard_analysis_2016). The PEP framework allowed us to numerically drive our theoretical proofs, and to assess the sharpness of our bounds.
This paper is organized as follows: Section 2 presents our main results in the convex setting, while Section 3 focuses on the strongly convex setting. We discuss the tightness of our results in Section 4, supported by numerical experiments. All technical results and extensions of our results are gathered in the appendices.
2 Convex Setting
In this section, we consider the smooth convex setting, which we formalize below. Note that the assumption could be relaxed, see the discussion in Section 5.1.
Assumption 2.1 (Convexity and smoothness).
Considering Problem (Finite-Sum), we assume that each function $f_{i}$ is convex and $L$-smooth for some $L\in(0,+\infty)$.
As mentioned in the introduction, our analysis is based on Lyapunov energies of the form (Lyapunov). The next lemma illustrates how a Lyapunov decrease leads to upper bounds for (SGD).
Lemma 2.2 (Bound from Lyapunov decrease).
Let Assumption 2.1 hold. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, that $\rho>0$, and, without loss of generality, that $a_{0}=1$. Then
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$ and $\bar{e}=\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$.
The next results are consequences of Lemma 2.2, after finding specific values of $\rho$, $(a_{k})$ and $(e_{k})$, aiming at the largest possible value for $\rho$. Most results will be presented with a uniform (and simpler) upper bound on $e_{k}$ for ease of presentation, and the more detailed constants can be found in Appendix C, whereto all proofs are postponed. The presentation of our results is split into three parts based on the normalized step-size $\gamma L$, namely Theorem 2.3 for short step-sizes $\gamma L\in(0,1)$, Theorem 2.4 for the optimal step-size $\gamma L=1$ and Theorem 2.7 for large step-sizes $\gamma L\in(1,2)$.
Theorem 2.3 (Convex case, short step-sizes).
Let Assumption 2.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,1)$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{L\|x_{0}-x_{*}\|^{2}}{2\gamma LT+2(1-\gamma L)}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}.$
The setting $\gamma L\in(0,1)$, usually considered short step-sizes in the GD literature, is the only one for which we have competitors in the literature. In (gower_sgd_2021, Theorem D.6) the authors obtain bounds of the same order as ours, albeit worse, for step-sizes limited to $\gamma L\in\left(0,\frac{1}{2}\right)$. They also obtain results by relaxing the convexity and smoothness assumptions, see more details in Appendix C.4. The authors in (garrigos_handbook_2024, Theorem 5.5) focus on the same assumptions as ours, and also obtain similar but worse bounds, for $\gamma L\in(0,\frac{1}{4}]$. We are not aware of any works for SGD without variance assumption providing bounds for the full range $\gamma L\in(0,1)$.
To the best of our knowledge, making a uniform variance assumption does not allow to obtain a bound with a better bias term than the one we obtained in Theorem 2.3. For instance (taylor_stochastic_2019, Theorem 6) obtain the same bias term as ours for short step-sizes, and liu_revisiting_2024 achieved a bias of $\frac{1}{\gamma L}$, although their results focuses on the last iterate.
Note that our result focuses on the Cesaro averages of the iterates $\bar{x}_{T}$, which is standard for SGD with no variance assumption. This is in contrast with the literature with uniformly bounded variance, for which it is possible to derive bounds on the last iterate $x_{T}$. The first result of this kind in the convex and smooth setting was established by bach_non-asymptotic_2011, with bounds recently improved in (taylor_stochastic_2019, Theorem 5) and (liu_revisiting_2024, Theorem 3.1). They obtain a variance term that scales with $T$ and $\log(T)$, respectively. Both were able to derive complexity bounds, with the latter offering a significantly better one over the former. It remains an open question to know whether or not it is possible to obtain bounds on the last iterate, similar to the one established in Theorem 2.3, without uniformly bounded variance. For the time being, this showcases a discrepancy between the uniformly bounded variance setting and ours.
We now turn to the case $\gamma L=1$, which we call the optimal step-size, with a slight abuse of terminology, in analogy with the optimal step-size for GD, and in view of the numerical results presented in Section 4.
Theorem 2.4 (Convex case, optimal step-size).
Let Assumption 2.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L=1$. Then, for every $T\geq 1$ and $\varepsilon\in(0,2)$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{(2-\varepsilon)\gamma T}+\frac{\gamma(2+\varepsilon)\sigma_{*}^{2}}{\varepsilon(2-\varepsilon)}.$
The case $\gamma L=1$, which is standard for GD, has surprisingly not yet seen any theoretical guarantees without bounded variance assumptions. As far as we know, this is the first upper bound provided in this setting. We point out the fact that Theorem 2.4 cannot be seen as the limit of Theorem 2.3 when $\gamma L\to 1$, because it would require to set $\varepsilon=0$, which results in an infinite variance term. Although it would seem like a flaw of our analysis that $\varepsilon=0$ is not feasible, we present evidence against this in Section 4.1. This singularity came as a surprise to us, and suggests that having a bias term of the order of $\tfrac{L}{2T}$ is impossible for SGD without any variance assumptions. This is a striking difference with GD, but also with SGD with bounded variance assumption, where no singularity holds at $\gamma L=1$, see for instance (taylor_stochastic_2019, Theorem 6).
Remark 2.5 (Stochastic proximal algorithm).
While the convergence rates for the deterministic proximal algorithm are well-studied, complexity results for its stochastic counterpart are unknown in a general setting. But a classic trick allows to rewrite the stochastic proximal algorithm as a particular instance of SGD, for a problem with no bounded variance, and with a step-size which must be $\gamma L=1$. Thanks to Theorem 2.4 we are now able to handle this optimal step-size and to derive the first complexity bound for this algorithm with no other assumption on the functions than convexity, avoiding, for instance, to impose finiteness and Lipschitzness as in davis_stochastic_2019. For more details, see Appendix H.
Interestingly, when $\sigma_{*}^{2}=0$ the variance term vanishes and allows us to take $\varepsilon\to 0$. This setting is known as the interpolation regime, because it is equivalent to ask that all the functions $f_{i}$ share a common minimizer, see (garrigos_handbook_2024, Lemma 4.17). It is a feature typically shared by overparametrized models.
Corollary 2.6 (Convex case with interpolation, optimal step-size).
Let Assumption 2.1 hold and assume that $\sigma_{*}^{2}=0$. Let $(x_{t})$ be generated by (SGD) with $\gamma L=1$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T},$
where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Theorem 2.7 (Convex case, large step-sizes).
Let Assumption 2.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(1,2)$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{\delta\|x_{0}-x_{*}\|^{2}}{2\gamma(2-\gamma L)T}+\frac{\gamma\bar{e}_{T}^{\prime}\sigma_{*}^{2}}{2(2-\gamma L)^{3}},$
where $\delta=1-(1-\gamma L)^{2T}\in(0,1)$, and $\bar{e}_{T}^{\prime}$ is exponential in $T$ (see its definition in Appendix C).
In this result, we obtain a bound where the variance term increases exponentially with $T$. This means that, for large step-sizes, the variance term will quickly dominate the bias term, preventing the algorithm from making significant progress. Of course, one might argue that this comment is not relevant because it is based on an upper bound, and that the proof of this upper bound could be improved. However, we provide empirical evidence in Appendix F which supports the claim that our bound is tight, and that the exponential increase in the variance term cannot be avoided.
This highlights a second discrepancy between the uniformly bounded variance setting and ours. In the large step-size regime, we are unable to provide a uniform bound of the variance term with respect to the time horizon $T$. This contrasts with the result in (taylor_stochastic_2019, Theorem 6), which, under the assumption of uniformly bounded variance, establishes a variance term that remains bounded with respect to $T$, for step-sizes $\gamma L\in(0,\tfrac{1+\sqrt{5}}{2})$.
3 Strongly Convex Setting
We now focus on the strongly convex setting. As in the convex setting, the below assumption may be relaxed, as we discuss further in Section 5.1.
Assumption 3.1 (Strong convexity and smoothness).
Considering Problem (Finite-Sum), we assume that each function $f_{i}$ is $\mu$-strongly convex and $L$-smooth for some $\mu,L\in(0,+\infty)$.
Our analysis is still based on the Lyapunov energy introduced in (Lyapunov). The following lemma explains how a decrease in the energy yields an upper bound for the iterates of (SGD).
Lemma 3.2 (Bound from Lyapunov decrease).
Let Assumption 3.1 hold. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\ldots,T-1$, $\rho\geq 0$, and that, without loss of generality, $a_{T}=1$. Then

### v2

# 3Université Paris Cité and Sorbonne Université, CNRS, Laboratoire de Probabilités, Statistique et Modélisation, F-75013 Paris, France

@@ Abstract
The non-asymptotic analysis of Stochastic Gradient Descent (SGD) typically yields bounds that decompose into a bias term and a variance term. In this work, we focus on the bias component and study the extent to which SGD can match the optimal convergence behavior of deterministic gradient descent. Assuming only (strong) convexity and smoothness of the objective, we derive new bounds that are bias-optimal, in the sense that the bias term coincides with the worst-case rate of gradient descent. Our results hold for the full range of constant step-sizes $\gamma L\in(0,2)$, including critical and large step-size regimes that were previously unexplored without additional variance assumptions. The bounds are obtained through the construction of a simple Lyapunov energy whose monotonicity yields sharp convergence guarantees. To design the parameters of this energy, we employ the Performance Estimation Problem framework, which we also use to provide numerical evidence for the optimality of the associated variance terms.
@@ 1 Introduction
Given a family $\{f_{i}\}_{i\in\mathcal{I}}$ of convex and smooth real-valued functions $f_{i}\colon\mathcal{H}\to\mathbb{R}$, we consider the problem given by
$\min_{x\in\mathcal{H}}f(x),\quad\text{where }f\coloneqq\mathbb{E}_{{i\sim\mathcal{D}}}\mathopen{}\left[{f_{i}}\right]\mathclose{},$ ($\mathbb{E}$-minimization)
where $\mathcal{D}$ is a probability distribution over $\mathcal{I}$. We assume $\operatorname{argmin}(f)$ to be nonempty. When the indexing set $\mathcal{I}$ is finite, note that we recover the typical finite sum minimization problem.
The Stochastic Gradient Descent (SGD) algorithm (robbins_stochastic_1951, ) is one of the most popular approaches for solving Problem ($\mathbb{E}$-minimization), and is widely used for large-scale machine learning and stochastic optimization problems. In its standard version, at each iteration, SGD performs a gradient step using one of the functions, which is chosen at random. More precisely, given the iterate $x_{t}$, it picks $i_{t}$ independently and identically distributed according to $\mathcal{D}$, and computes $x_{t+1}$ via
@@
where $\gamma>0$ is the step-size. Intuitively, SGD is able to progress towards a solution of the problem because it follows a direction that, on average, is an unbiased estimator for $\nabla f$. The simplicity and efficiency of SGD make it the go-to algorithm for training deep neural networks and other models on massive datasets, where computing the entire gradient $\nabla f$ is intractable or too expensive. While we keep our focus on vanilla SGD, all the results in this paper also apply to mini-batching and non-uniform sampling. This is a simple consequence of the the arguments laid out in gower_sgd_2019 , which we also provide in Appendix C for completeness and for the readers convenience.
Complexity of SGD. Quantitative results on the behavior of SGD consist in upper bounds for a metric $\Delta(T)$, which is to be interpreted as an optimality gap at iteration $T$. Typical choices for such metric include the expected solution gap $\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}$, the expected function value gap $\mathbb{E}\mathopen{}\left[{f(x_{T})-\inf f}\right]\mathclose{}$, the expected average function value gap $\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}$, or the expected ergodic function value gap $\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\inf f}\right]\mathclose{}$, where $\bar{x}_{T}$ is the Cesàro average of the iterates. In most known results garrigos_handbook_2024 , these upper bounds are expressed as the sum of two terms, taking the form
@@
where, typically, the bias term vanishes when $T\to+\infty$ and the variance term can be made arbitrarily small by taking a small enough step-size. For strongly convex and smooth problems, the metric $\Delta(T)$ will typically be the expected solution gap. For convex smooth problems, most proofs construct a bound on the expected average function value gap, and then derive a bound on the expected ergodic function value gap through Jensen’s inequality.
This paper discusses and addresses the following question: what are the best non-asymptotic bounds for SGD? In recent years, remarkable efforts have been made to answer this question for various deterministic optimization methods. These results include the optimal convergence rates of gradient descent (drori_performance_2014, ; taylor_smooth_2017, ; rotaru_exact_2024, ), proximal gradient descent (taylor_exact_2018, ), gradient descent with line-search (de_klerk_worst-case_2017, ), to name a few. In contrast, obtaining an optimal non-asymptotic complexity of stochastic algorithms is significantly more challenging and the optimal complexity of SGD remains nowadays unknown. In this work, we establish new and improved bounds for SGD, that are provably optimal in some sense. These bounds take the form
@@
where $x_{*}$ is a solution of the problem and $\sigma^{2}_{*}=\mathbb{V}[\nabla f_{i}(x_{*})]$ is the solution gradient variance constant (see more details in Section 2). We note that in the deterministic setting, where SGD reduces to GD and where $\sigma^{2}_{*}=0$, such bound reduces to $\Delta(T)\leq\text{Bias}(T)\cdot\|x_{0}-x_{*}\|^{2}$. We will therefore say that a bound is bias-optimal if the bias term matches the optimal convergence rate of GD. In particular, a bias-optimal bound would guarantee that SGD attains an unimprovable convergence rate of the metric $\Delta(T)$ in the interpolation regime (that is when $\sigma^{2}_{*}=0$). We note, however, that it is not clear whether general bounds for SGD always take the form of a gradient descent bound plus an additional variance term. In this work, we solely focus on bounds of the given form.
Contributions. In the convex case, we establish new and improved bounds in expectation for the average function value gap, which automatically convert into bounds for the more standard ergodic function value gap. Our bounds, obtained in Lemmas 4.1, 4.2 and 4.3, with simplified notation for the ease of presentation, are, for arbitrary $\varepsilon\in(0,2)$,
[…]
For short step-sizes $\gamma L\in(0,1)$, our bounds provide better constants than what was previously known. More interesting, we prove that our bound is bias-optimal for both the expected average function value gap and the expected ergodic function value gap. The setting $\gamma L\in[1,2)$ is more delicate to summarize. Our bounds are totally new, in the sense that this setting was never studied before. More importantly, they reveal and suggest non-trivial interactions between the bias and variance terms which, up to our knowledge, were never observed before. We now split our discussion between the critical and large step-sizes.
For large step-sizes $\gamma L\in(1,2)$, we prove that our bound is bias-optimal for the expected average function value gap. A striking feature of this bound is that the variance term has an impractical exponential grow with time, which is a phenomenon previously unobserved. We were not able to prove formally that this variance term cannot be improved, but our numerical experiments (see more details in the dedicated section) suggest that one cannot avoid a variance term diverging with time whilst keeping an optimal bias term. However, if we allow for a slightly worse bias term, hence deviating from the bias-optimal setting, this variance may be uniformly bounded with respect to $T$ (see Lemma 4.4 for more details). Of course, those considerations about the variance are irrelevant in the interpolation regime, where $\sigma_{*}^{2}=0$.
Regarding the critical step-size $\gamma L=1$, it is natural to conjecture that the optimal bias should be $\tfrac{L}{2T}$, which is the limiting optimal bias as $\gamma L\to 1$. This constant $\tfrac{L}{2T}$ plays a particular role, because it is also the infimum of the optimal biases for $\gamma L\in(0,2)$, and is also the best constant for GD. We prove that it is possible to obtain a bound for which the bias term is arbitrarily close to $\tfrac{L}{2T}$, but that it comes at the cost of a constant variance term which diverges. We moreover conjecture that it is not possible to derive a bound with a bias term equal to $\tfrac{L}{2T}$ when $\gamma L=1$, which we support by numerical evidence.
@@
We highlight that the study of the critical step-size $\gamma L=1$ in the convex smooth setting is novel. This in turn unlocks the study of the stochastic proximal algorithm, by viewing it as an instance of SGD applied to regularized functions. To the best of our knowledge, we establish the first complexity guarantees for this method in a general convex nonsmooth setting, considering possibly functions which do not have a full domain. This is presented in Section 6.
Let us now turn to the smooth strongly convex case, where each function $f_{i}$ is $L$-smooth and $\mu$-strongly convex. Our performance metric is the expected distance to the solution, namely $\Delta(T)=\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}$. Conceptually, our results are of the same nature as in the convex setting. Firstly, we obtain bounds for the full range of step-sizes $\gamma L\in(0,2)$, while the existing literature (bach_non-asymptotic_2011, ; needell_stochastic_2016, ; gower_sgd_2019, ) considered at best $\gamma L<1$. Extending the range of valid step-sizes allows us to consider the critical step-size $\gamma=\tfrac{2}{\mu+L}>1/L$, which is known to be optimal in the deterministic setting. Second, we show that our bounds are bias-optimal, except for the critical step-size for which we observe again a singularity. For every non-critical step-size we obtain ${\rm Bias}(T)=\phi^{2T}$, where $\phi$ defines the optimal constant for GD, namely
$\phi=\max\{1-\gamma\mu;\gamma L-1\}.$
For the critical step-size, we show that $\phi^{2T}$ can be approached with an arbitrary tolerance, with the cost of a variance growing to infinity. Supported by numerical results, we again conjecture that this optimal bias cannot be obtained.
@@
Proof Strategy. For each fixed time horizon $T$, we define a time-dependent random energy sequence
@@
where $\rho$, $(a_{t})$ and $(e_{t})$ are nonnegative parameters. The first term in the Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which decreases for deterministic gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. We chose to replace the standard term $t(f(x_{t})-\min f)$ with the sum of the past function gaps, which is of the same order in time. Up to our knowledge, this choice is new in the context of stochastic optimization, and we believe this choice to be critical to be able to derive bias-optimal bounds. The third term is a negative cumulated sum, introduced to compensate the fluctuations caused by the variance in the SGD algorithm, and hence to allow the Lyapunov energy to decrease.
@@
A decrease of $\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ along the iterations implies that $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, which translates into a bound for SGD (see Section 3.1 for more details). Our strategy is therefore to identify admissible Lyapunov parameters, which are those that make the energy decrease in expectation. Among those admissible parameters, our goal is to find the ones minimizing, whenever possible, the bias term in our bound.
To this end, we rely on the approach followed in taylor_stochastic_2019 (see also lessard_analysis_2016 ; fercoq_defining_2024 ; upadhyaya_automated_2025 ) to study Lyapunov potentials in a systematic fashion. It is based on the Performance Estimation Problem (PEP) methodology, which originally appeared in drori_performance_2014 , and was further developed in taylor_smooth_2017 . The PEP framework allows to find numerically admissible Lyapunov parameters, but also the ones minimizing the corresponding bias, by reformulating such problem into a semidefinite program. It helped us in two ways. First it helped us to navigate the space of admissible Lyapunov parameters and to obtain initial guesses for some parameter sequences, which eventually led us to our theoretical proofs. The reader curious about this process may check the proofs in the appendix for a more in-depth description. Second, it allowed us to numerically assess certain conjectures, such as the bias-optimality of our bounds, or the singularity of critical step-sizes.
Related Work. In early studies of SGD in the convex (smooth) framework, it was standard to make strong assumptions on the stochastic gradients $\nabla f_{i_{t}}(x_{t})$. The oldest, and possibly the best-known, are the uniform boundedness of the variance, and its variant, the uniform boundedness of the gradients, namely (robbins_stochastic_1951, ; nemirovski_robust_2009, ; rakhlin_making_2012, ; schmidt_fast_2013, )